OpenStudy (phebe):

MEDALSSS!!! 1. Which number makes this equation true? **v^2+10v+16=(v+8)(v+?) A-2 B-4 C-6 D-8

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} v^2&+10v&+16\\ &\uparrow &\uparrow \\ &8+2&8\cdot 2 \end{array}\) what do you think?

OpenStudy (phebe):

so its A right

OpenStudy (phebe):

20 for 10*2

OpenStudy (phebe):

so is it B-4

OpenStudy (phebe):

so it is B

OpenStudy (anonymous):

Jdoe gave a better response. What you should do is plug in the possible answers for the question mark, and FOIL the right side, to see if it looks like the left. Again, sorry about that.

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} v^2&+10v&+16\\ &\uparrow &\uparrow \\ &\square +\square &\square \cdot \square \end{array}\bf \implies (v+\square )(v+\square )\)

OpenStudy (phebe):

so its b-4 rit